OpenStudy (anonymous):

How do I solve this question? 9a + 7b = 30 8b + 5c = 11 -3a + 10c = 73 What do I eliminate first?

OpenStudy (anonymous):

And please give out a explanation of what ur doing :)

OpenStudy (jdoe0001):

hmmm, well... maybe I'm not looking ... much... but the elimination method didn't give me much

OpenStudy (jdoe0001):

the substitution method may work though

OpenStudy (jdoe0001):

are you supposed to use elimination?

OpenStudy (anonymous):

Sorry I just saw this I was doing another question

OpenStudy (anonymous):

I guess yea

OpenStudy (john_es):

I would elminate the third, is easier. Row1+3*Row3,

OpenStudy (anonymous):

I kind of did it and im not getting the 3rd point

OpenStudy (anonymous):

@john_es

OpenStudy (john_es):

Multiply the third row by 3, 9a + 7b = 30 8b + 5c = 11 -9a + 30c = 219 Sum the first and third row, 9a + 7b = 30 8b + 5c = 11 7b+30c = 249 Now multiply the third by -8 and the second by 7, sum, and substitute the result in the third row, 9a+7b=30 8b+5c=11 -205c=-1915 Solve the last equation, c=383/41 Do backsubstitution, a=279/41 b=-183/41

OpenStudy (anonymous):

Thanx :)